# Democratizing AI scientists using ToolUniverse

### v1

@@ Abstract
AI scientists are emerging computational systems that serve as collaborative partners in discovery. These systems remain difficult to build because they are bespoke, tied to rigid workflows, and lack shared environments that unify tools, data, and analyses into a common ecosystem. In omics, unified ecosystems have transformed research by enabling interoperability, reuse, and community-driven development; AI scientists require comparable infrastructure. We present ToolUniverse, an ecosystem for building AI scientists from any language or reasoning model, whether open or closed. ToolUniverse standardizes how AI scientists identify and call tools, integrating more than 600 machine learning models, datasets, APIs, and scientific packages for data analysis, knowledge retrieval, and experimental design. It automatically refines tool interfaces for correct use by AI scientists, creates new tools from natural language descriptions, iteratively optimizes tool specifications, and composes tools into agentic workflows. In a case study of hypercholesterolemia, ToolUniverse was used to create an AI scientist to identify a potent analog of a drug with favorable predicted properties. The open-source ToolUniverse is available at https://aiscientist.tools.
@@
AI scientists hold promise as computational systems that can reason, experiment, and collaborate in discovery [1, 2]. Yet most require one-off implementation, remain constrained by rigid workflows, and lack shared environments for reuse and growth [3, 2]. In contrast, fields such as omics have advanced through ecosystems that integrate tools and standardized analyses into common platforms [4, 5, 6]. Comparable infrastructure is needed to support AI scientists. Here we introduce ToolUniverse, a general ecosystem for constructing AI scientists at scale by combining large language models (LLMs), AI agents, and large reasoning models (LRMs) with an ecosystem of scientific tools, including machine learning models, datasets, and APIs (Figure 1a). Because scientific progress requires interaction with the world rather than reasoning alone [7, 1], ToolUniverse provides interactive environments where models can invoke tools, run experiments, and incorporate real-world feedback, transforming internal cognition into concrete research actions. By abstracting these capabilities behind a unified interface, ToolUniverse provides an ecosystem that wraps around any AI model, i.e., LLM, AI agent, or LRM, and enables the user to create and refine their own entirely custom AI research assistant (i.e., AI scientist) without the need for additional training or finetuning.
@@
ToolUniverse integrates more than 600 tools that span machine learning models, agents, software utilities, robotics, databases, and APIs (Figure 1b). To overcome incompatibilities between tools, it implements a tool specification schema that standardizes tool definitions and enables consistent inference across models. Analogous to the role of HTTP in regulating internet communication (Figure 1c), ToolUniverse establishes an AI-tool interaction protocol that governs how AI scientists issue tool requests and receive results. The protocol implements two operations (Figure 1d) in ToolUniverse: Find Tool, which maps natural-language descriptions to tool specifications, and Call Tool, which executes a selected tool with arguments and returns structured results such as text, embeddings, or JSON. This unified interface allows AI scientists to integrate tools at scale without bespoke setup.
@@
ToolUniverse is continually expanded with new tools. New tools can be registered locally or remotely and integrated without additional configuration [8]. Tools with complex dependencies or restricted access are supported through remote connections. ToolUniverse creates new tools from natural language descriptions, optimizes tool specifications iteratively, and composes interoperable tools. It chains tools for sequential or parallel execution so that AI scientists built with ToolUniverse can orchestrate workflows in a self-directed manner.
@@
At the core of ToolUniverse is the AI-tool interaction protocol that defines tools and standardizes how AI scientist systems interact with them. This protocol comprises three elements: a specification schema, an interaction schema, and communication methods. The specification schema provides a common format for describing each tool’s function, parameters, and outputs, enabling any client (whether an LLM, agent, or human user) to invoke tools without knowledge of their internal implementation. The interaction schema defines requests in a uniform way, encoding function calls with tool names and arguments. This standardization makes diverse tools interchangeable, whether they are local functions, remote machine learning models, or laboratory instruments. Communication protocols manage execution. Local operations run directly in Python, while remote serving uses the Model Context Protocol (MCP) to transmit requests across networks [8]. These abstractions make heterogeneous tools accessible through a consistent and extensible interface.
@@
Tool Finder (Figure 1e). The tool finder identifies relevant tools from over 600 resources using three complementary strategies: a keyword search for rapid retrieval, an LLM-based in-context search for semantic understanding, and an embedding search for scalable similarity matching. By combining speed, precision, and depth, AI scientist systems can locate tools to execute user tasks.
@@
Tool Discover (Figure 1i). Tool Discover generates new tools from natural language descriptions. It synthesizes formal specifications, produces executable implementations, validates outputs, and iteratively refines quality. This process ensures that newly created tools are maintainable and can be integrated into ToolUniverse.
@@
ToolUniverse integrates with LLMs, reasoning models, and agents to create customized AI scientist systems capable of planning, selecting tools, running experiments, and refining hypotheses. Setup requires only three steps: installing ToolUniverse, connecting it to a chosen model, and providing the model with a scientific problem. Once configured, the AI scientist can identify relevant tools, execute them, interpret results, and request human feedback when needed.
Three main approaches illustrate how ToolUniverse supports both general-purpose and specialized AI scientists. First, LLMs such as Claude or GPT can be equipped with tool access through simple in-context instructions [11] or lightweight configuration, enabling them to invoke and chain tools during reasoning (Figure 2a). Second, agentic systems such as Gemini CLI can directly use ToolUniverse’s MCP server and Tool Finder to identify and call tools, requiring minimal user setup (Figure 2b). Third, specialized agents, such as TxAgent for medical research [10] and Virtual Lab for nanobody design [12], can integrate with ToolUniverse at inference and during training. In the latter case, reinforcement learning strengthens their ability to navigate complex, task-specific environments [13, 14]. (Figure 2c).
@@
Through this screening process, the AI scientist identifies pravastatin, a drug with lower off-target effects than lovastatin [20]. In addition to reproducing established findings, ToolUniverse identifies a new candidate molecule (CHEMBL2347006/CHEMBL3970138) predicted to bind with higher affinity to HMG-CoA reductase, exhibit reduced blood-brain barrier penetrance, and display improved oral bioavailability and metabolic stability compared with lovastatin. Subsequent evidence confirmed that this compound had been patented for cardiovascular indications. The predicted binding affinity, binding likelihood, and BBB penetrance for all candidates are reported in Table LABEL:table:chembl_values. Using ToolUniverse, the AI scientist system selected, chained, and executed domain-specific tools to progress from hypothesis generation to candidate validation, while incorporating human feedback where needed. The workflow produced two candidates: an FDA-approved statin (pravastatin) and a patented small molecule. Both address lovastatin’s off-target effects due to activity outside the liver, one of its primary limitations. The patented small molecule also has higher predicted binding affinity than lovastatin, suggesting greater potency.
@@ Discussion
ToolUniverse moves beyond bespoke AI agents by providing an ecosystem for constructing AI scientist systems. It provides easy-to-use tools and operations, integrates new tools, composes tools into workflows, and incorporates automatic tool discovery and optimization. These capabilities extend any language model, agent, or reasoning model with research functionality tailored to the user: models can retrieve data or run existing scripts, and also identify relevant analysis tools, execute them with user inputs, combine outputs into multi-step workflows, and iteratively refine or even generate new tools when gaps are encountered. ToolUniverse provides an environment where reasoning models can carry out end-to-end research tasks, from hypothesis generation to experimental validation. It is applicable across scientific domains, and its integration of human-in-the-loop tools provides safeguards against erroneous outputs. For example, an AI scientist might propose several candidate compounds, and a human expert could validate tissue-specific expression before advancing them. Looking ahead, ToolUniverse could support multi-agent collaborations, benchmarking against human research groups, and integration with laboratory systems.
@@
ToolUniverse goes beyond orchestration frameworks such as GPT Agent [21], Gemini CLI [22], Claude Code [23], Qwen Code [24], LangChain [25], Haystack [26], Autogen [27] or CAMEL [28], and beyond communication protocols like MCP [8], which standardize how models access data resources, browsers, and tools. These systems are designed to route queries or connect agents with pre-defined tools, but they do not support the end-to-end lifecycle of creating, refining, and integrating tools into scientific workflows. In contrast, ToolUniverse can create and optimize tools and integrate them into agentic workflows beyond providing static tool registries. It introduces Tool Discover, which generates new tools from natural language descriptions and transforms them into ready-to-use components, and Tool Optimizer, which iteratively improves tool specifications through feedback from users and other AI systems. These capabilities enable ToolUniverse to manage and compose heterogeneous resources (Figure 1b) into new tools and expand the scientific toolkit.
We established the reliability of tools in ToolUniverse through a multi-step process that combines test example sampling, expert human review, and automated optimization to verify correctness and usability. ToolUniverse also undergoes regular maintenance with a structured bug reporting system that sustains long-term reliability. By drawing on tools from trusted, peer-reviewed, and regulatory sources, the ecosystem upholds scientific rigor.
A useful parallel for ToolUniverse is the rise of shared platforms in omics, where standardization reshaped how tools, data, and researchers interacted. These platforms created a foundation for reuse, reproducibility, and community-driven growth, driving major advances across the field. ToolUniverse serves a similar role for AI scientists by making AI agents broadly accessible and reducing the barriers to their use in research.
@@
Data and code availability. The project page and web service of ToolUniverse are at https://aiscientist.tools. The code, docs, and demos of ToolUniverse are at https://github.com/mims-harvard/ToolUniverse. The Python package of ToolUniverse is at https://pypi.org/project/tooluniverse. The case study of ToolUniverse is at https://zitniklab.hms.harvard.edu/ToolUniverse/tutorials/tooluniverse_case_study.html.
@@
Figure 1: a) ToolUniverse: an ecosystem for creating AI scientists. General-purpose LLMs, reasoning models, and agents are connected to the ToolUniverse ecosystem of over 600 scientific tools, enabling autonomous research workflows across domains. b) Overview of the diverse tool categories supported in ToolUniverse, including ML models, agents, domain knowledge, experimentation, scientific packages, automation, human feedback, workflows, datasets, APIs, embedding stores, visualization, and retrieval. c-d) Just as HTTP standardizes client–server communication, ToolUniverse provides an interaction protocol that regulates how AI models issue tool requests and receive responses. Core operations for AI scientists interacting with ToolUniverse: Find Tool (mapping natural language to tool specifications) and Call Tool (executing a tool and returning structured outputs). e) Tool finder: component for identifying relevant tools using three strategies: keyword search, LLM-based in-context search, and embedding-based similarity search. f) Tool caller: execution engine that validates inputs, dynamically loads tools, dispatches calls through ToolUniverse.run() or MCP, and returns structured outputs to clients. g) Tool manager: environment for integrating local and remote tools. Local tools are registered via JSON specifications and decorators, while remote tools are added through MCP for privacy- or dependency-constrained setups. h) Tool composer: environment for chaining multiple tools into composite workflows. Supports sequential, parallel, and feedback-driven orchestration of heterogeneous tools. i) Tool discover: multi-agent system for generating new tools from natural language requirements. Combines specification synthesis, automated code generation, validation, and iterative refinement to create production-ready tools. j) Tool optimizer: multi-agent system that iteratively refines tool specifications to improve clarity, accuracy, and usability. Integrates test generation, execution analysis, and feedback-driven improvement.
@@
Figure 2: a) Building AI scientists by connecting LLMs (such as Claude) with specialized tools. By selecting the right tools for each task and providing clear instructions, these LLM-powered AI scientists can effectively carry out scientific tasks. b) Building AI scientists with multi-round tool use and reasoning by connecting ToolUniverse with AI agents (such as Gemini CLI). These agents leverage tool-finding capabilities to select relevant tools, reason, and use ToolUniverse iteratively to tackle complex scientific tasks. c) ToolUniverse be applied to therapeutic discovery for hypercholesterolemia by connecting with Gemini CLI as an AI scientist. This example illustrates how ToolUniverse enables an AI scientist to select, chain, and use domain-specific tools, moving from hypothesis generation to candidate validation, while maintaining the flexibility to incorporate human feedback when appropriate.

### v2

@@ Abstract
AI scientists are emerging computational systems that serve as collaborative partners in discovery. These systems remain difficult to build because they are bespoke, tied to rigid workflows, and lack shared environments that unify tools, data, and analyses into a common ecosystem. In genomics, unified ecosystems have transformed research by enabling interoperability, reuse, and community-driven development; AI scientists require comparable infrastructure. We present ToolUniverse, an ecosystem for building AI scientists from any language or reasoning model across open- and closed-weight models. ToolUniverse standardizes how AI scientists identify and call tools by providing more than 600 machine learning models, datasets, APIs, and scientific packages for data analysis, knowledge retrieval, and experimental design. It automatically refines tool interfaces for correct use by AI scientists, generates new tools from natural language descriptions, iteratively optimizes tool specifications, and composes tools into agentic workflows. In a case study of hypercholesterolemia, ToolUniverse was used to create an AI scientist to identify a potent analog of a drug with favorable predicted properties. The open-source ToolUniverse is available at https://aiscientist.tools.
@@
AI scientists hold promise as computational systems that can reason, experiment, and collaborate in discovery [1]. Yet most require one-off implementation, remain constrained by rigid workflows, and lack shared environments for reuse and growth [2, 3]. In contrast, fields such as genomics have advanced through ecosystems that integrate tools and standardized analyses [4, 5, 6, 7, 8]. Comparable infrastructure is needed to support AI scientists [9]. Here we introduce ToolUniverse, a general ecosystem for constructing AI scientists at scale by combining large language models (LLMs), AI agents, and large reasoning models (LRMs) with scientific tools, including machine learning models, bioinformatics workflows, retrieval systems, remote and local datasets, scientific packages for data analysis, visualization and lab automation, and human-in-the-loop and safety tools (Figure 1a). As scientific research requires interaction with the real world rather than text-based reasoning alone [10, 1], ToolUniverse implements an environment where AI models can invoke tools, run experiments, and incorporate feedback from human experts or computational models. This environment wraps around a user-specified AI model, such as LLM, AI agent, or LRM, to create a customized AI research assistant (i.e., an AI scientist) tailored to user instruction without additional model training or fine-tuning.
ToolUniverse includes more than 600 tools spanning machine learning models, agents, software packages, robotics, datasets, and remote endpoints such as APIs and knowledge bases (Figure 1b). AI models face challenges in selecting and using these tools effectively, yet being able to use tools is essential for autonomous scientific research [2]. Scientific tools differ in purpose and complexity, from molecular docking predictors to literature retrieval APIs, and often require distinct input formats and parameters. Many are multimodal, processing genomic sequences, protein structures, microscopy images, or clinical text, while others perform Python-based modeling or statistical analyses that depend on specific runtime environments. Some operate locally, whereas others run remotely through APIs or laboratory automation systems. Overlapping and redundant functionalities between tools mean that each tool must be precisely described so that AI scientists can select and invoke the right ones for a given task.
@@
To address these compatibility issues of AI scientists using tools and making the large toolbox in ToolUniverse available for effective use by AI scientists, analogous to the role of HTTP in regulating internet communication (Figure 1c), AI-tool interaction protocol in ToolUniverse governs how AI scientists issue tool requests and receive results of executing the tools. The protocol implements two operations (Figure 1d) in ToolUniverse: Find Tool, which maps natural-language descriptions of tools to tool specifications that are understood by the underlying AI model that powers the AI scientist, and Call Tool, which executes a selected tool with arguments and returns the result, such as text, embeddings, or JSON objects.
@@
ToolUniverse is continually expanded with new tools. New tools can be registered locally or remotely and integrated without additional configuration [11]. Tools with complex dependencies or restricted access are supported through remote connections. As we describe below, ToolUniverse creates new tools from natural language descriptions, optimizes tool specifications iteratively, and composes interoperable tools. It chains tools for sequential or parallel execution so that AI scientists built with ToolUniverse can orchestrate workflows in a self-directed manner.
@@
At the core of ToolUniverse is the AI-tool interaction protocol that defines tools and standardizes how AI scientist systems interact with them. This protocol comprises three elements: a specification schema, an interaction schema, and communication methods. The specification schema provides a common format for describing each tool’s function, parameters, and outputs, enabling any client (whether an LLM, agent, or human user) to invoke tools without knowledge of their internal implementation. The interaction schema defines requests in a uniform way, encoding function calls with tool names and arguments. This standardization makes diverse tools interchangeable, whether they are local functions, remote machine learning models, or laboratory instruments. Communication protocols manage execution. Local operations run directly in Python, while remote serving uses the Model Context Protocol (MCP) [11] to transmit requests across networks. These abstractions make heterogeneous tools accessible to AI scientists through a consistent and extensible protocol.
@@
Tool Finder (Figure 1e). The Tool Finder identifies relevant tools from more than 600 tools using three strategies: keyword search for rapid retrieval, LLM-based in-context search for semantic understanding, and embedding search for scalable similarity matching. By flexibly identifying tools relevant to a user-specified task, ToolUniverse allows AI scientists built within the ecosystem to efficiently locate and execute task-relevant tools.
@@
Tool Discoverer (Figure 1i). Tool Discoverer generates new tools from natural language descriptions. It synthesizes formal specifications, produces executable implementations, validates outputs, and iteratively refines quality. This component ensures that newly created tools are maintainable and can be integrated into ToolUniverse.
@@
ToolUniverse integrates with LLMs, reasoning models, and agents to create customized AI scientists capable of planning, selecting tools, running experiments, and refining hypotheses. Setup requires only three steps: installing ToolUniverse, connecting it to a user-chosen AI model, and providing the model with a scientific problem. Once configured, the AI scientist can identify relevant tools, execute them, interpret results, and request human feedback when needed.
Three main approaches illustrate how ToolUniverse supports both general-purpose and specialized AI scientists. First, LLMs such as Claude or GPT can be equipped with tool access through simple in-context instructions [14] or a lightweight configuration, enabling them to invoke and chain tools during reasoning (Figure 2a). Second, agentic systems such as Gemini CLI can directly use ToolUniverse’s MCP server and Tool Finder to identify and call tools, requiring minimal user setup (Figure 2b). Third, specialized agents such as TxAgent for medical research [13], Virtual Lab for nanobody design [15], and GeneAgent for gene set function discovery [16] can use tools from ToolUniverse during training to improve tool use and reasoning [17] or during inference to expand their capabilities [18]. In the former case, reinforcement learning enables effective multi-tool collaborative reasoning [19, 20] (Figure 2c).
@@
Through this screening process, the AI scientist identifies pravastatin, a drug with lower off-target effects than lovastatin [26]. In addition to reproducing established findings, ToolUniverse identifies a new candidate molecule (CHEMBL2347006/CHEMBL3970138) predicted to bind with higher affinity to HMG-CoA reductase, exhibit reduced blood-brain barrier penetrance, and display improved oral bioavailability and metabolic stability compared with lovastatin. Subsequent evidence confirmed that this compound had been patented for cardiovascular indications. The predicted binding affinity, binding likelihood, and BBB penetrance for all candidates are reported in Table LABEL:table:chembl_values. Using ToolUniverse, the AI scientist system selected, chained, and executed domain-specific tools to progress from hypothesis generation to candidate validation, while incorporating human feedback where needed. The workflow produced two candidates: an FDA-approved statin (pravastatin) and a patented small molecule. Both address lovastatin’s off-target effects due to activity outside the liver, one of its primary safety issues. The patented small molecule also has higher predicted binding affinity than lovastatin, suggesting greater potency.
@@ Discussion
ToolUniverse moves beyond bespoke AI agents by providing an ecosystem for constructing AI scientists. It integrates tools and operations into workflows and incorporates automatic tool discovery and optimization. These capabilities extend any language model, agent, or reasoning model with research functionality tailored to the user: models can retrieve data or run existing scripts, and also identify relevant analysis tools, execute them with user inputs, combine outputs into multi-step workflows, and iteratively refine or even generate new tools when gaps are encountered. It is applicable across scientific domains, and its integration of human-in-the-loop tools provides safeguards against erroneous outputs. For example, an AI scientist might propose several candidate compounds, and a human expert could validate tissue-specific expression before advancing them. Looking ahead, ToolUniverse will be used to build AI scientists for discovery and serve as a testbed to evaluate their integration with laboratory systems, potential for safety and biosecurity risks [27], and adherence to safeguards [28] and governance specifications [29].
@@
ToolUniverse goes beyond orchestration frameworks such as GPT Agent [30], Gemini CLI [31], Claude Code [32], Qwen Code [33], LangChain [34], Haystack [35], Autogen [36] or CAMEL [37], and communication protocols like MCP [11] and AutoTools [38], which standardize how models access data resources, browsers, and tools. These systems are designed to route queries or connect agents with pre-defined tools, but they do not support the end-to-end lifecycle of creating, refining, and integrating tools into scientific workflows. In contrast, ToolUniverse can create and optimize tools and integrate them into agentic workflows beyond providing static tool registries. It introduces Tool Discoverer, which generates new tools from natural language descriptions and transforms them into ready-to-use components, and Tool Optimizer, which iteratively improves tool specifications through feedback from users and other AI systems. These capabilities enable ToolUniverse to manage and compose existing tools (Figure 1b) into new tools.
To ensure that tools can be effectively used by AI models to build AI scientists, ToolUniverse employs a multi-step validation process combining tool-specific test cases for each tool added to ToolUniverse, human expert reviews of input-output traces, and automated optimization of tool specifications to verify correctness and usability by AI scientists. Regular maintenance and a bug-reporting system further sustain long-term reliability.
A close parallel to ToolUniverse is the emergence of computational platforms in genomics, which created shared infrastructure for reuse, reproducibility, and open science innovation. ToolUniverse provides a comparable foundation for AI scientists, lowering barriers to their development and use in research.
@@
Data and code availability. The project page and web service of ToolUniverse are at https://aiscientist.tools. The code, docs, and demos of ToolUniverse are at https://github.com/mims-harvard/ToolUniverse. The Python package of ToolUniverse is at https://pypi.org/project/tooluniverse. The therapeutic case study is at https://zitniklab.hms.harvard.edu/ToolUniverse/tutorials/tooluniverse_case_study.html.
@@
Figure 1: a) ToolUniverse is an ecosystem for building AI scientists. General-purpose LLMs, reasoning models, and agents connect to the ToolUniverse ecosystem of more than 600 scientific tools to perform autonomous research workflows across domains. Illustrated are four open- and closed-weight AI models: GPT, Claude, Gemini, and DeepSeek. b) Overview of the tool categories in ToolUniverse, including ML models, agents, domain knowledge, experimentation, scientific packages, automation, human feedback, workflows, datasets, APIs, embedding stores, visualization, and retrieval. c-d) Like HTTP standardizes client–server communication, ToolUniverse defines an interaction protocol that governs how AI models issue tool requests and receive responses. Core operations for AI scientists interacting with ToolUniverse are Find Tool (map natural language to tool specifications) and Call Tool (execute tools and return tool outputs). e) Tool Finder identifies relevant tools using keyword search, LLM-based in-context search, and embedding-based similarity search. f) Tool Caller validates inputs, dynamically loads tools, dispatches calls through ToolUniverse.run() or MCP, and returns tool outputs. g) Tool Manager integrates local and remote tools. It registers local tools via JSON specifications and decorators and adds remote tools through MCP for privacy- or dependency-constrained setups. h) Tool Composer chains multiple tools into composite workflows. It supports sequential, parallel, and feedback-driven orchestration of heterogeneous tools. i) Tool Discoverer is a multi-agent system that generates new tools from natural language requirements. It performs specification synthesis, automated code generation, validation, and iterative refinement to produce production-ready tools. j) Tool Optimizer is a multi-agent system that iteratively refines tool specifications to improve clarity, accuracy, and usability. It combines test generation, execution analysis, and feedback-driven refinement. ML, Machine Learning; LLM, Large Language Model; HTTP, Hypertext Transfer Protocol; MCP, Model Context Protocol; RAG, Retrieval-Augmented Generation; API, Application Programming Interface; ELN/LIMS, Electronic Lab Notebook and Laboratory Information Management System; SDK, Software Development Kit.
@@
Figure 2: a) Building AI scientists by connecting LLMs (for example, Claude) with specialized tools. The AI scientist selects appropriate tools for each task, follows clear instructions, and executes tools from ToolUniverse. b) Building AI scientists with multi-round tool use and reasoning by connecting ToolUniverse with AI agents (for example, Gemini CLI). These agents identify relevant tools, reason across steps, and iteratively use ToolUniverse to solve complex, multi-stage scientific problems. c) Applying ToolUniverse to therapeutic discovery for hypercholesterolemia by connecting it with Gemini CLI to create an AI scientist. The analysis proceeds from target identification to compound screening, property optimization, and patent verification. The AI scientist identifies HMG-CoA reductase as a therapeutic target, screens existing and novel compounds using DrugBank, ChEMBL, and predictive ML tools such as Boltz-2 and ADMET-AI, and evaluates candidates through patent-mining tools to confirm novelty and prior art. This case study shows how AI scientists chain and execute domain-specific tools, incorporate human feedback, and validate therapeutic hypotheses through multi-step reasoning. Tools from ToolUniverse used in this illustration of the AI scientist include local datasets (“drugbank get drug name and description by target name”, “drugbank get drug name and description by indication”, “drugbank get pharmacology by drug name or drugbank id”), remote datasets and API tools (“OpenTargets get disease ID description by name”, “OpenTargets get associated targets by disease efoID”, “OpenTargets get target tractability by ensemblID”, “EuropePMC search articles”, “HPA search genes by query”, “HPA get rna expression in specific tissues”, “HPA get comprehensive gene details by ensembl ID”, “ChEMBL search similar molecules”, “PubChem get CID by SMILES”, “PubChem get associated patents by CID”, “get webpage text from URL”), feedback and safety (Consult human expert tool), and ML models (Boltz-2 and ADMET-AI tools). CLI, Command Line Interface; HPA, Human Protein Atlas; ADMET, Absorption, Distribution, Metabolism, Excretion, and Toxicity; CID, Compound Identifier.